1. the length of a rectangle is greater than twice its breadth by 4cm. If the perimeter of the rectangle is 26cm. Find its length and breadth.
2. The sum of the present ages of Ram and Shaym is 36 yrs. Ram is elder than Shaym by 2 rs. Find Their present ages.

Answer by simranb(63)

1.Let the breadth be x
Then the length=2x+4
Perimeter=26cm
perimeter=2(l+b)
so
2(l+b)=26
2(2x+4+x)=26
2(3x+4)=26
6x+8=26
6x=26-8
6x=18
x=18/6
x=3
then 2x+4=2*3+4=>6+4=>10
So length=10cm
Breadth=3cm
Your second problem i solved already.Just check it out.
Hope this helps!